OpenStudy (anonymous):

Find the area between each curve and the x-axis for the given interval. y=4x^2+2x from x=1 to x=4

OpenStudy (anonymous):

\[\int_1^4(4x^2+2x)dx\] is a start

geerky42 (geerky42):

yeah. remember: \[\large \int cx^n dx = c\int x^n dx =\dfrac{cx^{n+1}}{n+1}+C\]